OpenStudy (anonymous):

Find an equation for the nth term of the sequence. -2, -8, -32, -128, ...

OpenStudy (anonymous):

-2 = (-2)^1 -8 = (-2)^3 -32 = (-2)^5 -128 = (-2)^7 we can extrapolate, such that the nth term of the sequence = (-2)^n

OpenStudy (anonymous):

so (an = 4 • -2n) is right @tridian

OpenStudy (anonymous):

@heyheyhelpme

OpenStudy (anonymous):

What!?!?!!?!?!!?

OpenStudy (anonymous):

I don't know the answer!